OpenStudy (anonymous):

Solve for the roots of x in each of the equations below. x^4 - 81 = 0 x^4 + 10x^2 + 25 = 0 x^4 - x^2 - 6 = 0

OpenStudy (anonymous):

what number squared is 81?

OpenStudy (anonymous):

3, -3, 3i, -3i

OpenStudy (anonymous):

@rkopen

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

For question 1. \(\Large\color{black}{ \bf hint.~~~81=3^4 }\) For question 2. Let x^2=a, and complete the square. For question 3. Let x^2=a and factor.

OpenStudy (anonymous):

for 1 i have 3,-3,3i,-3i

OpenStudy (anonymous):

For 2 I have, 5,-5,5i,-5i

OpenStudy (anonymous):

hello

OpenStudy (solomonzelman):

I was out for a little, sorry about that. 1) Correct 2) \(\normalsize\color{black}{ \ x^4 + 10x^2 + 25 = 0 ~~~~~~~~~~let~~x^2=a }\) \(\normalsize\color{black}{ \ a^2 + 10a + 25 = 0 }\) \(\normalsize\color{black}{ \ (a+5)^2 = 0 }\) \(\normalsize\color{black}{ \ a+5 = 0 }\) \(\normalsize\color{black}{ \ a = -5 ~~~~~~~~~x^2=-5 }\) \(\normalsize\color{black}{ \ Hence,~~~~~ x = ±5i }\) -5 is not one of the x solutions.

OpenStudy (solomonzelman):

I mean +5 is not... -5 is correct. So 2) real root -5 imaginary roots +5i, -5i

OpenStudy (solomonzelman):

For number 3, \(\normalsize\color{blue}{ \ x^4 - x^2 - 6 = 0~~~~~~~~~~~~~let~~x^2=a}\) \(\normalsize\color{blue}{ \ a^2 - a - 6 = 0~~~~~~~~~~~~~~~factor,}\) \(\normalsize\color{blue}{ \ (a-3)(a+2) = 0}\) \(\normalsize\color{blue}{ \ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ Hence,~~~~ a=3,-2}\) \(\normalsize\color{blue}{ \ a=3,-2~~~~thus~}\) \(\normalsize\color{blue}{ \ x^2=3~~~~~~and~~~~~~x^2=-3}\) \(\normalsize\color{blue}{ \ x=± ~3~~~~~~and~~~~~~x=±~3i}\)

OpenStudy (anonymous):

3,-2,-3,+/-3,+/-3i or 3,-3, +/-3, +/-3i

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \ x^2=3~~~~~~and~~~~~~x^2=-2}\) \(\normalsize\color{blue}{ \ x=± \sqrt{3}~~~~~~and~~~~~~x=±~2i}\)

OpenStudy (solomonzelman):

That is what I meant.